OpenStudy (anonymous):

Convert the given function into a logarithmic function f(x)=2.01e^(0.14x)

OpenStudy (anonymous):

maybe the question is "find the inverse" ?

OpenStudy (anonymous):

i got this far .. y=2.01e^(0.14x) y/2.01= 2.01e^(0.14x)/2.01 In y/2.01=in e^(0.14x)

OpenStudy (anonymous):

its not ..

OpenStudy (anonymous):

but that is what you are finding right?

OpenStudy (anonymous):

no, im converting f(x)=2.01e^(0.14x) into log form

OpenStudy (anonymous):

?

OpenStudy (anonymous):

\[f(x)=2.01e^{0.14x}\] \[y=2.01e^{0.14x}\] \[\frac{y}{2.01}=e^{0.14x}\] \[0.14x=\ln(\frac{y}{2.01})\] \[x=\frac{1}{0.14}\ln(\frac{y}{2.01})\] but that is a completely different function, it is the inverse

OpenStudy (anonymous):

Convert the given function into a logarithmic function (hint: think BNE BEN). Check out an example. f(x) = 106.16e(-0.73x) y = 106.16e(-0.73x) y over one hundred six point 16 = e(-0.73x) In y over one hundred six point 16 = In e(-0.73x) In y - In 106.16 = In e-0.73x In y = In e-0.73x+ 4.66 In y = -0.73x+4.66

OpenStudy (anonymous):

seee its not inverse , omg this is so confusing ..

OpenStudy (anonymous):

ok your function is \(f(x)\) and then you replace \(f\) by \(y\) so when you change it you do not have \(f\) anymore but some other function

OpenStudy (anonymous):

okay so can you show me how to change y=2.01e^(0.14x) into a logarithmic function ?

OpenStudy (anonymous):

i guess you have the log of your original function i hate to keep repeating myself, but here \(y\) is an exponential function, it is not a logarithmic one you can find \(\ln(y)\) if you like

OpenStudy (anonymous):

we can run through the steps like in the previous example i think

OpenStudy (anonymous):

is In the same thing as y but in log form instead of exponential form ? .. idek what In is ,,

OpenStudy (anonymous):

start with \[y=2.01e^{0.14x}\] divide by 2.01 get \[\frac{y}{2.01}=e^{0.14x}\] take the log get \[\ln(\frac{y}{2.01})=0.14x\] rewrite the left hand side as \[\ln(y)-\ln(20.1)=0.14x\] use a calculator to find \(\ln(2.01)=.698\) to to get \[\ln(y)-.698=0.14x\] and finally write \[\ln(y)=0.14x+.698\]

OpenStudy (anonymous):

but at the risk of repeating myself, this is not \(y\) on the left, but it is \(\ln(y)\) not the same function, but rather the log of the original function that is what your example found

OpenStudy (anonymous):

oh, okay . i get it now! can you help me with one more thing pleaseeee ?

OpenStudy (anonymous):

sure

OpenStudy (anonymous):

okay well its like an activity and i had to use coins, and i did 7 trials and counted how many landed on heads and how many were on tails each time , i had to graph them by using the trial number as x. and the number of ones in the heads column as the y . i got (1,2) (2,3) (3,3) (4,6) (5,2) (6,2) (7,5) and then i had to Convert the given function into a logarithmic function which we already did :) but now ...i have to describe what the logarithmic graph will look like...which idfk how to do and Write a half-life function based on your data.

OpenStudy (anonymous):

wow you really lost me on this one but maybe we can figure it out

OpenStudy (anonymous):

what does this mean (1,2) (2,3) (3,3) (4,6) (5,2) (6,2) (7,5)

OpenStudy (anonymous):

i mean if you have a log function maybe you can find the half life, but i am rather confused

OpenStudy (anonymous):

those are the points i had to graph, the x intercepts at the trials 1,2,3,4,5,6,7 & the y intercepts are the amout of time the quarter landed on heads for each trial

OpenStudy (anonymous):

the x intercepts are *

OpenStudy (anonymous):

still lost how can a coin land heads more than once on a trial?

OpenStudy (anonymous):

i used 8 coins

OpenStudy (anonymous):

oh, that makes more sense

OpenStudy (anonymous):

so how do i describe what the logarithmic graph will look like...and Write a half-life function based on that data.

OpenStudy (anonymous):

i don't even know how to get a log function out of this try reposting the problem with only that question you will get a better answer than i can give

OpenStudy (anonymous):

what log function did you get?

OpenStudy (anonymous):

In(y)=0.14x+.698